OpenStudy (anonymous):

Solve the system by substitution. -x - y - z =-8 -4x + 4y + 5z = 7 2x + 2z = 4

OpenStudy (campbell_st):

well take the last equation and make z the subject... what you you get..?

OpenStudy (anonymous):

z=2-x

OpenStudy (anonymous):

Can you tell me if I am heading in the right direction? Substitute z into equation 2. -4x+4y+5(2-x)=7 -4x+4y+10-x=7 -4x+4y-x=-3 -3x+4y=-3

OpenStudy (campbell_st):

yes and you also need to make the substitution into the 1st equation...

OpenStudy (anonymous):

Wait, so where do I go from where I left off? Do I solve for x or y?

OpenStudy (campbell_st):

what you will end up with is 2 equations in 2 unknowns.... if z = 2 - x then in equation 1 you will have -x - y -(2 -x) = -8 so -x -y -2 + x = 8 simplifying, x will go -y = 10 so y = -10 now make 2 substitutions into the 2nd equation -4x +4(-10) +5(2 - x) = 7 which simplifies to -9x -30 = 7 solve for x. then substitute into the last equation to find z

OpenStudy (anonymous):

Thank you. I appreciate your help. :)

OpenStudy (anonymous):

I got -4.1 for x? Is it okay that it is a decimal?

OpenStudy (anonymous):

There was an error in your work... You had positive 8 when it should have been negative 8. I figured it out though. Thanks!

OpenStudy (campbell_st):

well I think x is closer to -4.2 repeater... this is an add question, in that the solutions aren't integers... and its about the 5 times I've seen it in the past fortnight. but ignoring the answers I hope the method makes sense... thats the important part